OpenStudy (anonymous):

write a polynomial function with real coefficients in standard form whose zeros include -2, 3 and 4-5i

OpenStudy (anonymous):

\[(x+2)(x-3)(x-(4+5i))(x-(4-5i))\] is start

OpenStudy (anonymous):

multiplying the first two terms is easy, you get \[(x+2)(x-3)=x^2-x-6\]

OpenStudy (anonymous):

it is the second two that is somewhat annoying, but you have a couple says to do it

OpenStudy (anonymous):

one way is to work backwards if a zero is \(4+5i\) then you can write \[x=4+5i\] \[x-4=5i\] \[(x-4)^2=-25\] \[x^2-8x+16=-25\] \[x^2-8x+41=0\] is your polynomial

OpenStudy (anonymous):

another way is to remember that if \(a+bi\) is the zero of a quadratic polynomial, the quadratic is \[x^2-2ax+a^2+b^2\] and in your case \(a=4,b=5\)